The weather has taken a turn for the worse at LCO. It is currently 1.8C and blowing at 32.7 mph. And getting colder.
Speaking of cold, we installed the new liquid cooling system in the W-Unit today.
The other big accomplishment in world of the NAS today was Laird installing the Telecentric Lens. This almost flat lens lets us work off-axis, away from the guide star.
We also got some time in the dome, and did a little work on the Clay telescope. The vane-end trays are installed, and the leak preventing solenoid valve was plumbed and tested.
This hawk was hanging out over the observatory at lunch today.
Tonight’s pretty pictures are from Alan Uomoto, who spent some time messing with his camera’s settings today.
Days without a motherboard failure: 2